Find 4 consecutive odd integers with a sum of 184.
Answer #1

Find 4 consecutive odd integers with a sum of 184.

1st: 2x-3 2nd: 2x-1 3rd: 2x+1 4th: 2x+3

Equation: Adding the above you get 8x = 184 x = 23

1st: 2x-3 = 43 2nd: 45 3rd: 47 4th: 49

answered by: J.R.
Answer #2
Let the first odd integer be f Then the 2nd, 3rd, &4th are f + 2, f + 4, &f + 6, respectively Since their sum is 184, then we get: f + (f + 2) + (f + 4) + (f + 6) = 184 4f + 12 = 184 4f = 172 {{{f = 172/4 = 43}}} Since the 1st number is 43, it follows that the four consecutive odd numbers are {{{highlight_green(43)}}},{{{highlight_green(45)}}},{{{highlight_green(47)}}} &{{{highlight_green(49)}}}.
answered by: kmk